# Phase retrieval via media diversity

### v1

@@ Abstract
This work studies phase retrieval for wave fields, aiming to recover the phase of an incoming wave from multi-plane intensity measurements behind different types of linear and nonlinear media. We show that unique phase retrieval can be achieved by utilizing intensity data produced by multiple media. This uniqueness does not require prescribed boundary conditions for the phase in the incidence plane, in contrast to existing phase retrieval methods based on the transport of intensity equation. Moreover, the uniqueness proofs lead to explicit phase reconstruction algorithms. Numerical simulations are presented to validate the theory.
Key words. phase retrieval, media diversity, wave propagation, nonlinear media, nonlinear Schrödinger equations, inverse problems, transport of intensity equation
@@ 1 Introduction
Besides the advances in numerical reconstruction algorithms [9, 15, 24, 47, 48], a significant amount of efforts in the area has been (necessarily) spent on designing measurement strategies to enrich the diversity of the phase information encoded in the intensity data from which phase data are extracted. Examples of such strategies include, but are not limited to, for instance, measuring intensity data at multiple transverse planes (i.e. multi-plane measurement) [19, 20], exploiting phase-amplitude interactions due to nonlinear propagation [29, 30, 35], enhancing partial coherency via structured multi-frequency, multi-probe, or multi-illumination measurements [21, 25, 26, 37, 38], and many more [2, 12, 14, 39].
@@
where $0\leq Z_{1}<Z_{2}\leq L_{z}$, while $\mathcal{P}$ and $\mathcal{D}$ denote respectively the parameter (i.e. medium) and data spaces of the problem.
@@
In Appendix A, we present some simple examples to show that one can not expect uniqueness in phase retrieval with intensity data from a single medium, even if the medium is nonlinear and one can measure intensity data at infinitely many transverse planes. We will see, however, that the quadratic Schrödinger model (3) actually allows us to regain uniqueness with data from multiple media.
@@ 2 Recovering phase gradient with linear media diversity
Throughout the rest of the paper, we need to take various forms of derivatives of the intensity of the wave field $u$. Here we make some general assumptions on the parameters of the problem so that one can reasonably believe that such differentiations are well-defined. We refer to [10, 6, 22, 23] and references therein for more detailed mathematical discussions on the regularity of solutions to the equations discussed here.
@@
Such assumptions, together with appropriate boundary conditions when $\Omega$ is a bounded domain (for instance, homogeneous Dirichlet boundary conditions), are often sufficient to ensure the smoothness of solutions to (1), (3), and (5) in $z$ and $\mathbf{x}$ that we need in the calculations in this section and the next. Interested readers are referred to  [10] and references therein for rigorous mathematical treatment on the subject.
@@
This gives immediately
@@
To make sense of the complicated right-hand side, we recognize that after multiplying through by $\kappa$ it is of the form
@@ 3 Recovering phase with nonlinear media diversity
The second equation in particular informs us that
@@
The same calculation can be conducted to show that this technique also works for nonlinear media with the term $\beta u^{2}$ replaced with $\beta u^{3}$. However, the technique also fails to work for linear media with term $\beta u$. Also, the technique will not work with media governed by the nonlinearity given by the Gross-Pitaevskii model (5); see Remark 2.5.
@@ 4.1 Reconstruction from $\kappa$ diversity
at every $\mathbf{x}\in\Omega$. For $m=1,2,\dots,N_{m}$, the notation $u_{m}$ denotes the solution to (1) with $\kappa=\kappa_{m}$. Given measured intensity data, we can evaluate the right-hand side by a finite difference scheme, and then solve the system, in least-square sense when $N_{m}\geq d+2$, to find $F_{1}$, $F_{2}$ and $F_{3}$. Meanwhile, from (15), we have that
@@ 4.2 Reconstruction from $\beta$ diversity
In the uniqueness theory, we need intensity data from at least $N_{m}\geq d+2$ well-selected (to make the $\mathcal{K}$ matrix invertible) $\kappa$ media to reconstruct the phase gradient $\nabla\varphi$ and data from at least $N_{m}\geq d+4$ well-selected (to ensure the invertibility of the matrix $\mathcal{B}$) $\beta$ media to uniquely reconstruct the phase $\varphi$. In practice, it would always be better to obtain data from more media, say $N_{m}\gg d+4$, for better stability. In such cases, the linear systems (27) and (31) can be solved in the least-squares manner.
@@
The uniqueness proofs in Section 2 and Section 3, as well as the reconstruction method derived from them, are essentially based on the so-called transport of intensity equations [48]; see (7) and (20). Standard ways of using this technique require boundary conditions for the phase, which is often impractical. Media diversity allows us to avoid such a requirement in our method. This is an essential difference between what we have here and those in the literature.
@@ 5 Numerical experiments
These $\kappa$ profiles are selected to ensure the matrix $\mathcal{K}$ in (6) is full-rank when a sufficiently large number of $\kappa$ is used as $\mathcal{K}$ being invertible is a necessary condition for the reconstruction process. Two typical reconstructions are provided in Figure 2. Shown are the true phase gradient (top row), the reconstruction with data from the first four media $\kappa_{1}\sim\kappa_{4}$ (middle row), and the reconstruction with data from all five media $\kappa_{1}\sim\kappa_{5}$.
@@ 6 Computational retrieval in general settings
Our primary parameter of interest is $f(\mathbf{x})=a(\mathbf{x})\exp(i\varphi_{\text{true}}(\mathbf{x}))$, and the amplitude of $f$, $a(\mathbf{x})$, is known to us. Therefore, the only parameter we aim to reconstruct is the phase $\varphi_{\text{true}}(\mathbf{x})$. Since our numerical experiments are also performed using synthetic data, we always know the true phase $\varphi_{\rm true}$. To quantify the error in the reconstruction $\varphi_{\rm reconstructed}$, we define its mean shift from the truth by
@@ 6.2 Multi-media data
Figure 7: Phase reconstructions for the quadratic nonlinear Schrödinger model (3) with single-plane, multiple media data. Shown are reconstructions by optimization algorithm with three different randomly selected initial guesses.
@@ 7 Concluding remarks
This work studies phase retrieval for wave fields using linear and nonlinear media diversity, aiming to recover the phase of an incoming wave from intensity measurement behind multiple media. While such a phase reconstruction is generally impossible, we show that it can be uniquely achieved when we have a well-selected group of media. Moreover, when this is achievable, direct (i.e. non-iterative) methods can be developed for the reconstruction of the phase from measured intensity data. While our derivation of the phase reconstruction method is based on the transport of intensity equation, our method does NOT require boundary conditions for the phase, making it dramatically different from similar methods in the literature that are based on the transport of intensity equations.
@@ A.2 Non-uniqueness for the GP equation
Similar nonuniqueness examples can be constructed for the GP model (5). Recall that we can only hope to reconstruct the initial condition $f$ up to a constant phase shift. Therefore, we henceforth identify initial conditions which differ by a constant phase and consider the problem of reconstructing the equivalence class of $f$ under this equivalence relation.
@@
To focus on the main idea, we set $\Omega=\mathbb{R}^{2}$ in order to not have to deal with boundary conditions. Let us define a subset $\mathcal{B}^{\prime}\subset\mathcal{B}$ as follows. We say that $\beta=\beta(\mathbf{x})\in\mathcal{B}$ is in $\mathcal{B}^{\prime}$ if: there exists a constant $c$ and a non-constant harmonic function $\psi\in\mathcal{C}^{2}(\overline{\Omega};\mathbb{R})$ such that
@@ B.1 Adjoint state gradient calculation
The same calculations can be performed to solve the phase retrieval problem for the Gross-Pitaevski equation. Indeed, we replace the quadratic nonlinear Schrödinger model in the optimization problem (39) with the Gross-Pitaevski equation with periodic boundary conditions:
@@ B.2 Domain discretization
Our simulation covers the domain $(0,L_{z})\times\Omega$ with $\Omega=(0,L_{x})\times(0,L_{y})$. We discretize $\Omega$ by $100\times 100$ cells of uniform size. The spatial grid sizes are $dx=dy=0.01$, and for numerical stability and resolution considerations, we choose $dz=1\times e^{-5}$. The total simulation time is fixed at $L_{z}=0.02$. This time frame allows the waves to travel properly within the domain without using too much computational resources.
@@ B.3 Finite difference discretizations of PDEs
The forward models (40) and (45), as well as the corresponding adjoint equations (43) and (46) are all discretized using a standard time-domain finite difference method. In particular, we use the fourth-order Runge Kutta method for the time variable.

### v2

@@ Abstract
This work studies phase retrieval in the Fresnel regime, aiming to recover the phase of an incoming wave from multi-plane intensity measurements behind different types of linear and nonlinear media. We show that unique phase retrieval can be achieved by utilizing intensity data produced by multiple media. This uniqueness does not require prescribed boundary conditions for the phase in the incidence plane, in contrast to existing phase retrieval methods based on the transport of intensity equation. Moreover, the uniqueness proofs lead to explicit phase reconstruction algorithms. Numerical simulations are presented to validate the theory.
Key words. phase retrieval, media diversity, Fresnel regime, wave propagation, nonlinear media, nonlinear Schrödinger equations, inverse problems, transport of intensity equation
@@ 1 Introduction
Besides the advances in numerical reconstruction algorithms [9, 15, 24, 47, 48], a significant amount of efforts in the area have been (necessarily) spent on designing measurement strategies to enrich the diversity of the phase information encoded in the intensity data from which phase data are extracted. Examples of such strategies include, but are not limited to, for instance, measuring intensity data at multiple transverse planes (i.e. multi-plane measurement) [19, 20], exploiting phase-amplitude interactions due to nonlinear propagation [29, 30, 35], enhancing partial coherency via structured multi-frequency, multi-probe, or multi-illumination measurements [21, 25, 26, 37, 38], and many more [2, 12, 14, 39].
@@
where $0\leq Z_{1}<Z_{2}\leq L_{z}$, while $\mathcal{P}$ and $\mathcal{D}$ denote respectively the parameter (i.e., medium) and data spaces of the problem.
@@
In Appendix A, we present some simple examples to show that one can not expect uniqueness in phase retrieval with intensity data from a single medium, even if the medium is nonlinear and one can measure intensity data at infinitely many transverse planes. We will see, however, that data from multiple media can potentially allow us to restore this uniqueness.
@@ 2 Recovering phase gradient with linear media diversity
Throughout the rest of the paper, we need to take various forms of derivatives of the intensity of the wave field $u$. Here, we make some general assumptions on the parameters of the problem so that one can reasonably believe that such differentiations are well-defined. We refer to [10, 6, 22, 23] and references therein for more detailed mathematical discussions on the regularity of solutions to the equations discussed here.
@@
Such assumptions, together with appropriate boundary conditions when $\Omega$ is a bounded domain (for instance, homogeneous Dirichlet boundary conditions), are often sufficient to ensure the smoothness of solutions to (1), (3), and (5) in $z$ and $\mathbf{x}$ that we need in the calculations in this section and the next. Interested readers are referred to  [10] and references therein for rigorous mathematical treatments of the subject.
@@
This immediately gives
@@
To make sense of the complicated right-hand side, we recognize that after multiplying through by $\kappa$, it is of the form
@@ 3 Recovering phase with nonlinear media diversity
The second equation, in particular, informs us that
@@
The same calculation can be conducted to show that this technique also works for nonlinear media with the term $\beta u^{2}$ replaced with $\beta u^{3}$. However, the technique also fails to work for linear media with the term $\beta u$. Also, the technique will not work with media governed by the nonlinearity given by the Gross-Pitaevskii model (5); see Remark 2.5.
@@ 4.1 Reconstruction from $\kappa$ diversity
at every $\mathbf{x}\in\Omega$. For $m=1,2,\dots,N_{m}$, the notation $u_{m}$ denotes the solution to (1) with $\kappa=\kappa_{m}$. Given measured intensity data, we can evaluate the right-hand side by a finite difference scheme, and then solve the system, in the least-squares sense when $N_{m}\geq d+2$, to find $F_{1}$, $F_{2}$, and $F_{3}$. Meanwhile, from (15), we have that
@@ 4.2 Reconstruction from $\beta$ diversity
In the uniqueness theory, we need intensity data from at least $N_{m}\geq d+2$ well-selected (to make the $\mathcal{K}$ matrix invertible) $\kappa$ media to reconstruct the phase gradient $\nabla\varphi$ and data from at least $N_{m}\geq d+4$ well-selected (to ensure the invertibility of the matrix $\mathcal{B}$) $\beta$ media to uniquely reconstruct the phase $\varphi$. Besides uniqueness consideration, one also needs to select media combinations such that the matrices $\mathcal{K}$ and $\mathcal{B}$ have good condition numbers. Therefore, it is better to obtain data from more media, say $N_{m}\gg d+4$, for better stability. In such cases, the linear systems (27) and (31) can be solved in the least-squares manner.
@@
The uniqueness proofs in Section 2 and Section 3, as well as the reconstruction method derived from them, are essentially based on the so-called transport of intensity equations [48]; see (7) and (20). Standard ways of using this technique require boundary conditions for the phase, which is often impractical. Media diversity allows us to avoid such a requirement in our method. This is an essential difference between what we have here and what is in the literature.
@@ 5 Numerical experiments
These $\kappa$ profiles are selected to ensure the matrix $\mathcal{K}$ in (6) is full-rank when a sufficiently large number of $\kappa$ is used, as $\mathcal{K}$ being invertible is a necessary condition for the reconstruction process. Two typical reconstructions are provided in Figure 2. Shown are the true phase gradient (top row), the reconstruction with data from the first four media $\kappa_{1}\sim\kappa_{4}$ (middle row), and the reconstruction with data from all five media $\kappa_{1}\sim\kappa_{5}$.
@@
There are several important aspects of the numerical implementation of the phase reconstruction algorithm.
First, the quality of the numerical reconstruction of the phase of the incoming wave remains on the same level for different intensity profiles, as long as $|f|\geq\varepsilon>0$ for some $\varepsilon$ much larger than the machine accuracy. For instance, we can have almost identical phase reconstructions as the ones in Figure 3 using either $|f(\mathbf{x})|=1$ or $|f(\mathbf{x})|=\exp(-\frac{(x-0.5)^{2}+(y-0.5)^{2}}{0.5})$.
Second, instead of directly reconstructing the phase, the method we have reconstructs the (rescaled) real and imaginary parts of $f$, that is, $\nu$ and $\omega$ in the representation $f=a(\nu+i\omega)$ (where $a=|f|$). Our method does not explicitly enforce the constraint $\nu^{2}+\omega^{2}=1$ in the process. However, we observe that this constraint is satisfied very well for the successful reconstructions we performed. The reconstructions usually give $\nu^{2}+\omega^{2}=1$ up to a few digits of accuracy.
Third, the reconstruction methods we proposed are local, meaning that we need to invert matrices $\mathcal{K}$ and $\mathcal{B}$ at each point of the computational domain. These are small matrices whose inversion is computationally cheap to obtain. As we mentioned before, in our implementations, we take the simplest sets of media parameters $\kappa$ and $\beta$ to ensure that the matrices $\mathcal{K}$ and $\mathcal{B}$ are invertible at every point in the computational domain.
@@ 6 Computational retrieval in general settings
Let us emphasize again that while our main interest is the phase of the incident wave, our reconstruction algorithm attempts to reconstruct the incident wave $f$ itself. In other words, if $f(\mathbf{x})=a(\mathbf{x})\exp(i\varphi(\mathbf{x}))$, our algorithm reconstruct both the amplitude $|f|$ and the phase $\varphi$. If we also measure the amplitude of $f$, $a(\mathbf{x})$, we can add $|f|=a$ as a constraint in the optimization process. In such a case, it would be more convenient to view $\Phi$ as a functional $\Psi$ of $\varphi$, i.e.,
$\Psi(\varphi)=\Phi(a\exp(i\varphi)).$ (34)
The Fréchet derivative of $\Psi$ can be computed using the Fréchet derivative of $\Psi$ along with the chain rule; more details can be found in Appendix B.
Since our numerical experiments are also performed using synthetic data, we always know the true phase $\varphi_{\rm true}$. To quantify the error in the reconstruction $\varphi_{\rm reconstructed}$, we define its mean shift from the truth by
@@ 6.2 Multi-media data
Figure 7: Phase reconstructions for the quadratic nonlinear Schrödinger model (3) with single-plane, multiple media data. Shown are reconstructions by an optimization algorithm with three different randomly selected initial guesses.
@@ 7 Concluding remarks
This work studies phase retrieval for wave fields using linear and nonlinear media diversity, aiming to recover the phase of an incoming wave from intensity measurements behind multiple media. While such a phase reconstruction is generally impossible, we show that it can be uniquely achieved when we have a well-selected group of media. Moreover, when this is achievable, direct (i.e. non-iterative) methods can be developed for the reconstruction of the phase from measured intensity data. While our derivation of the phase reconstruction method is based on the transport of intensity equation, our method does NOT require boundary conditions for the phase, making it dramatically different from similar methods in the literature that are based on the transport of intensity equations.
@@ A.2 Non-uniqueness for the GP equation
Similar nonuniqueness examples can be constructed for the GP model (5). Recall that we can only hope to reconstruct the initial condition $f$ up to a constant phase shift. Therefore, we henceforth identify initial conditions that differ by a constant phase and consider the problem of reconstructing the equivalence class of $f$ under this equivalence relation.
@@
To focus on the main idea, we set $\Omega=\mathbb{R}^{2}$ in order not to have to deal with boundary conditions. Let us define a subset $\mathcal{B}^{\prime}\subset\mathcal{B}$ as follows. We say that $\beta=\beta(\mathbf{x})\in\mathcal{B}$ is in $\mathcal{B}^{\prime}$ if: there exists a constant $c$ and a non-constant harmonic function $\psi\in\mathcal{C}^{2}(\overline{\Omega};\mathbb{R})$ such that
@@ B.1 Adjoint state gradient calculation
Additionally, recall from (34) that when the amplitude of $f$ is given (as say $a$), the problem reduces to minimizing $\Psi(\varphi)=\Phi(a\exp(i\varphi))$. In this case, the Fréchet derivative of $\Psi(\varphi)$ at $\varphi$ in the direction $\delta\varphi$ can be computed from the Fréchet derivative (45) of $\Phi$ and the chain rule:
$\Psi^{\prime}(\varphi)[\delta\varphi]$ $=\Phi^{\prime}(a\exp(i\varphi))[ia\exp(i\varphi)\delta\varphi]$
[…]
The same calculations can be performed to solve the phase retrieval problem for the Gross-Pitaevskii equation. Indeed, we replace the quadratic nonlinear Schrödinger model in the optimization problem (40) with the Gross-Pitaevski equation with periodic boundary conditions:
@@ B.2 Domain discretization
Our simulation covers the domain $(0,L_{z})\times\Omega$ with $\Omega=(0,L_{x})\times(0,L_{y})$. We discretize $\Omega$ by $100\times 100$ cells of uniform size. The spatial grid sizes are $dx=dy=0.01$, and for numerical stability and resolution considerations, we choose $dz=1\times e^{-5}$. The total simulation time is fixed at $L_{z}=0.02$. This time frame allows the waves to travel properly within the domain without using too many computational resources.
@@ B.3 Finite difference discretizations of PDEs
The forward models (41) and (46), as well as the corresponding adjoint equations (44) and (47) are all discretized using a standard time-domain finite difference method. In particular, we use the fourth-order Runge-Kutta method for the time variable.